# On admissibility in post-hoc hypothesis testing

### v2

@@ 4 Properties of admissible rules
One property which follows from Lemma 4.1 is that any admissible test must be decreasing in the loss. In other words, as the loss increases, a test should only become more conservative.
@@
Lemma 4.8 will show that if $E$ defines a $\Gamma$-admissible binary test family for any $\Gamma$ then $E$ is compatible. Such a general result does not hold in the randomized case, which can be seen by considering the e-values of Neyman-Pearson tests when the distributions are discrete. However, Lemma 5.10 proves that when $P$ and $Q$ are continuous and the randomized test $\delta_{E}$ is $\mathbf{C}$-admissible, then $E$ is compatible. Interestingly, the issue of compatibility also arose for Grünwald (2024), even when studying a different notion of admissibility. He called the condition “richness.”
@@ 4.1 Randomized test families
We will sometimes call (4.2) the canonical representation of $\delta$, or simply say that $\delta$ is canonical if written as such. Lemma 4.5 will be particularly useful in the upcoming theorems on admissibility, since it lets us assume that all test families we deal with are canonical.
@@ 4.2 Binary test families
Throughout this section we assume that $P$ and $Q$ are continuous. We begin with the analogue of Lemma 4.5 but for binary test families. The same representation was given recently in Ramdas and Wang (2025, Chapter 4), though stated as a definition, not as the unique form of any admissible binary test family.
@@ 5.1 $\mathbf{U}$-admissibility
The invocation of Condition (C1) in Theorem 5.1 prohibits us from translating the results into the standard Neyman-Pearson setting. That is, Observation 2.5 does not apply. Indeed, for a singleton $\mathcal{B}=\{b\}$, (C1) implies that $L_{b}(1,0)=0$, so type-II risk loses its meaning and does not translate into the power of a test in the traditional paradigm.
@@
Intriguingly, Theorem 5.1 and Corollary 5.3 recover the result of Grünwald (2024) up to some assumptions, even though he studied a weaker form of admissibility. In particular, Theorem 1 of Grünwald (2024) essentially shows that any G-admissible rule is defined by a sharp e-variable. Roughly speaking, he considers a test family to be G-admissible if it is not dominated with probability 1 by another test. (We use G-admissibility to refer to Grünwald.)
@@ 5.2 Improving tests via Rao–Blackwellization
These examples might make us wonder whether all “good” e-variables for post-hoc hypothesis testing for point hypotheses are functions of the likelihood ratio. In general, of course, there are e-variables which are not functions of $\Lambda$. Indeed, $E=\text{d}R/\text{d}P$ for any distribution $R$ is a (sharp) e-variable. However, unless $R=Q$, such an e-variable is not log-optimal (Shafer, 2021). We might therefore expect that it can be improved as a post-hoc test.
Here we show that Rao-Blackwellization (Rao, 1945; Blackwell, 1947) can weakly improve the type-II risk of any test family based on an e-variable, with respect to a particular set of adversaries. Recall that the Rao-Blackwell theorem says that the expected loss of an estimator can be weakly improved (with respect to a convex loss) by conditioning on a sufficient statistic. See Wasserman (2004, Theorem 9.42) for a modern statement of the result.
@@
It is worth mentioning that by the same logic, we can show that $\mathbb{E}_{Q}[\log S_{T}]\geq\mathbb{E}_{Q}[\log E]$ or indeed, $\mathbb{E}_{Q}[f(S_{T})]\geq\mathbb{E}_{Q}[f(E)]$ for any nonnegative, concave function $f$. That is, Rao-Blackwellization leads to an e-variable that never has worse performance under $Q$, where performance is measured by some concave loss function. We state this as a theorem with a self-contained proof since it may be of independent interest.
Theorem 5.7.
Let $T$ be a sufficient statistic for testing $P$ vs $Q$. Let $E$ an e-variable for $P$ and define $S_{T}=\mathbb{E}_{P}[E|T]$. Then $S_{T}$ is an e-variable and $\mathbb{E}_{Q}[f(S_{T})]\geq\mathbb{E}_{Q}[f(E)]$ for any concave function $f$. In particular, $\mathbb{E}_{Q}[\log(S_{T})]\geq\mathbb{E}_{Q}[\log(E)]$.
Proof.
By concavity, $\mathbb{E}_{P}[f(E)|\sigma(T)]\leq f(\mathbb{E}_{P}[E|\sigma(T)])=f(S)$. As above, the likelihood ratio $\Lambda=dQ/dP$ is minimally sufficient for testing $P$ vs $Q$, so it is $\sigma(T)$-measurable. Therefore,
$\mathbb{E}_{Q}[f(S)]$ $=\mathbb{E}_{P}[\Lambda f(S)]\geq\mathbb{E}_{P}[\Lambda\mathbb{E}_{P}[f(E)|\sigma(T)]]$
$=\mathbb{E}_{P}[\mathbb{E}_{P}[\Lambda f(E)|\sigma(T)]]=\mathbb{E}_{P}[\Lambda f(E)]=\mathbb{E}_{Q}[f(E)],$
as desired. ∎
Note that $S_{T}(X)$ in (5.1) can be written as $S_{T}(X)=g(T(X))$ for some function $g$. Therefore, taking $T=\Lambda$, a consequence of Proposition 5.6 is that for appropriate adversarial classes $\Gamma$, any $\Gamma$-admissible test family based on an e-variable is a function of the likelihood ratio.
Corollary 5.8.
Let $E$ be an e-variable for $P$ and let $\delta_{E}$ be $\Gamma$-admissible for any $\Gamma\subseteq\Gamma_{\sigma(\Lambda)}$. Then $\delta_{E}$ is a function of the likelihood ratio.
@@ 5.3 $\mathbf{C}$-admissibility
We begin by proving that any randomized $\mathbf{C}$-admissible test must be increasing in the likelihood ratio. The result is, of course, satisfyingly intuitive—a larger likelihood ratio indicates indicates more evidence against the null. A test that rejects the null at $\Lambda(X)$ but fails to reject when $\Lambda(Y)>\Lambda(X)$ is doing so despite increased odds in favor of the alternative. That such a result does not hold for richer classes of adversaries $\Gamma$ is yet another reason to consider $\mathbf{C}$. Indeed, tests which don’t satisfy this property are flaunting the law of likelihood (Royall, 2017).
@@
The proof may be found in Appendix A.13. The idea is simple: Any test which is not increasing in $\Lambda$ can be perturbed such that it becomes strictly better. The details are intricate, however, requiring a careful analysis and strictly more edge cases than is healthy.
@@
It is worth noting that the properties given in Lemmas 5.9 and 5.10 are both on display for the likelihood ratio test. For continuous $P$ and $Q$, its associated e-value is $E^{\textsc{np}}(X)=L_{b^{*}}(0,1)$ if $\Lambda(X)\geq\kappa(b^{*})$ and 0 otherwise (see Section 3), which is clearly compatible with $\mathcal{L}$. However, in the discrete case, the e-value obeys $E^{\textsc{np}}(X)=L_{b^{*}}(0,1)(1+\gamma)\mathbf{1}\left\{\Lambda(X)\geq\kappa(b^{*})\right\}$, which may not belong to $\mathcal{L}$. Hence, the assumption of continuity in Lemma 5.10 cannot be removed.
We leave it as an open question whether the conditions specified by Lemmas 5.9 and 5.10 (in addition, of course, to the canonical representation in Lemma 4.5) are also sufficient to specify $\mathbf{C}$-admissible tests. We expect the answer is no. We do, however, conjecture that in addition to our extensions of the likelihood ratio test, mixtures of likelihood ratio tests are $\mathbf{C}$-admissible, subject to compatibility. More formally, if $\phi^{\textsc{np}}_{k}$ is the post-hoc likelihood ratio test on loss $b_{k}$ (i.e., the test defined by (3.2)) then we suspect that the test defined by the (sharp) e-variable
@@ 5.3.1 Binary test families
is $\mathbf{C}$-admissible. Indeed, $\Lambda$ is sharp, compatible with $\mathbb{R}_{\geq 0}$ and obviously increasing in itself. This test is satisfying in that it accords well with calls to report the likelihood ratio instead of the p-value (Royall, 1986, 2000; Berger and Sellke, 1987; Perneger, 2001). This nicely bridges the Fisherian and Neyman-Pearson perspectives on hypothesis testing: If a study summarizes its findings via $\Lambda(X)$ as a measure of evidence, downstream analysts can reject at whatever level they like, maintaining type-I risk safety.
@@ 6 Summary
Different families $\Gamma$ result in different classes of admissible tests. For point hypotheses, when $\Gamma$ is all mappings from the data to losses we show that all admissible tests correspond to sharp e-variables. When $\Gamma$ is the set of constant mappings and the tests are binary, we give a classification of all admissible tests which recovers the Neyman-Pearson lemma (Neyman and Pearson, 1933).
This work suggests several open questions. For one, in the case of randomized tests, the set of necessary and sufficient conditions for $\mathbf{C}$-admissibility do not match. Can we close the gap and give a classification of all admissible tests in this case? As discussed in Section 5.3, we expect that mixtures of likelihood ratio tests are admissible but we have been unable to prove it. Second, while we have focused mainly on point hypotheses, hypothesis testing in practice typically involves composite hypotheses. While some of our results extend straightforwardly to composite settings, [^2] In particular, Lemmas 4.1, 4.2, and 4.3 can be extended to composite settings if handled correctly. For instance, a composite version of Lemma 4.1 would posit that $\phi(X,b)\geq\delta(X,b)$ $P$-almost surely for all $P\in\mathcal{P}$ and that there exists some $Q\in\mathcal{Q}$ such that $Q(\phi(X,b)>\delta(X,b))>0$. The representations in Lemmas 4.5 and 4.7 likewise hold in composite settings. can we obtain composite versions of our main classification theorems? Third, we have adopted the typical hypothesis testing setup of two actions: either sustaining or rejecting the null. Grünwald (2024), however, studies multiple actions. How must our results be modified in this case?
@@ Acknowledgments
We thank Ruodo Wang for pointers towards measurable selection theorems. BC was supported in part by the CWI internship program and in part by NSERC-PGSD, grant no. 567944. BC and AR acknowledge support from NSF grants IIS-2229881 and DMS-2310718. PG was co-funded by the European Union (ERC Advanced Grant FLEX 101142168).
@@ Appendix B On the definition of admissibility
First let us discuss the notion of admissibility used by Grünwald (2024), which we henceforth refer to as G-admissibility. We say a test family $\delta$ is G-inadmissible if there exists a type-I risk safe test family $\phi$ such that, for all $b\in\mathcal{B}$ and all $Q\in\mathcal{Q}$,
@@
In other words, $\phi$ has loss at most that of $\delta$ with $Q$-probability 1 on all losses, and has lower loss than $\delta$ with positive probability. If no such $\phi$ exists then we say that $\delta$ is G-admissible. We note that Grünwald (2024) defines G-admissibility in terms of type-I losses instead of type-II, after preprocessing the losses to remove any such that $L_{b}(1,a)>L_{b}(1,a)$ and $L_{b}(0,a)>L_{b}(0,a)$ for some $a$, which can be done without loss of generality.
@@ B.1 Type-I risk as the supremum over adversaries
where $\Gamma_{\Omega}$ represents the projection of a set in $\Omega\times\mathcal{B}$ onto $\Omega$. The set $A_{c}$ is Borel (i.e., lies in $\Sigma_{\mathcal{X}}\otimes\mathcal{B}$) by the measurability of $L_{b}(0,1)\delta(X(\omega),b)$. The projection of a Borel set on a product space is analytic by Suslin’s theorem, hence $\{\omega:g(\omega)>c\}$ is analytic (Kechris, 2012, Exercise 14.3). Moreover, every analytic set is universally measurable (Kechris, 2012, Theorem 21.10), meaning it is measurable for any finite Borel measure on $\Omega$, and $P$ is finite and Borel by assumption. Finally, if the super-level sets of a real-valued function are measurable, then the function itself measurable. We conclude that $g$ is $\mathcal{F}$-measurable, so the left hand side of (2.3) exists. Now, for any $X$ and $B$, $L_{B(X)}(0,\delta(X,B(X)))\leq\sup_{b}L_{b}(0,\delta(X,b))$. Taking expectations and then taking the supremum over all maps $B$, we see that the right hand side of (2.3) is at most the left hand side.
@@
which is $\mathcal{F}$-measurable since $(\omega,b)\mapsto L_{b}(0,\delta(X(\omega),b))$ is measurable. The Kuratowski–Ryll-Nardzewski measurable selection theorem (Kuratowski and Ryll-Nardzewski, 1965) (see also Kechris 2012, Theorem 12.13) guarantees the existence of a measurable selector $B^{*}_{\epsilon}(X(\omega))$ such that $B_{\epsilon}(X(\omega))\in\Xi_{\epsilon}(\omega)$, hence $L_{B^{*}_{\epsilon}(X)}(0,\delta(X(\omega),B_{\epsilon}^{*}(X))+\epsilon>\sup_{b}L_{b}(0,\delta(X,b))$. Taking expectations and letting $\epsilon\downarrow 0$ shows that the left hand side of (2.3) is at most the right, which proves the result. ∎
@@
There exist pathological choices of $\{L_{b}\}_{b\in\mathcal{B}}$ which ensure that $\sup_{b}L_{b}(0,\delta(X,b))$ is not measurable, in which case Definition 2.3 breaks down and Lemma B.1 doesn’t hold. Instead of making various measurability assumptions as we do in this work, one could circumvent this problem by following Grünwald (2024) and calling $\delta$ type-I risk safe if there exists some measurable function $\xi:\mathcal{X}\to\mathbb{R}_{\geq 0}$ such that

### v3

@@ 4 Properties of admissible rules
One property which follows from Lemma 4.1 is that any admissible test must be decreasing in the loss. In other words, as the loss increases, a test should only become more conservative. This fact was also noted by koning2023markov, though not related to admissibility.
@@
Lemma 4.8 will show that if $E$ defines a $\Gamma$-admissible binary test family for any $\Gamma$ then $E$ is compatible. Such a general result does not hold in the randomized case, which can be seen by considering the e-values of Neyman-Pearson tests when the distributions are discrete. However, Lemma 5.8 proves that when $P$ and $Q$ are continuous and the randomized test $\delta_{E}$ is $\mathbf{C}$-admissible, then $E$ is compatible. Interestingly, the issue of compatibility also arose for grunwald2024beyond, even when studying a different notion of admissibility. He called the condition “richness.”
@@ 4.1 Randomized test families
We will sometimes call (4.2) the canonical representation of $\delta$, or simply say that $\delta$ is canonical if written as such. Lemma 4.5 will be particularly useful in the upcoming theorems on admissibility, since it lets us assume that all test families we deal with are canonical. We note that koning2023markov shows that the test in (4.2) is type-I risk safe.
@@ 4.2 Binary test families
Throughout this section we assume that $P$ and $Q$ are continuous. We begin with the analogue of Lemma 4.5 but for binary test families. The same representation was given recently in ramdas2024hypothesis, though stated as a definition, not as the unique form of any admissible binary test family. Let us note that koning2023markov proves a very similar result. In particular, after translating his result to our setting, he shows that given an e-variable $E$, we can obtain a type-I risk safe test via $(X,b)\mapsto\mathbf{1}\left\{L_{b}(0,1)\leq E(X)\right\}$. Conversely, given a type-I risk safe test, then $E_{\phi}=\sup\{L_{b}(0,1)\leq E(X)\}$ is an e-value. We show that, starting with $\phi$, one can write $\phi(X,b)=\mathbf{1}\left\{L_{b}(0,1)\leq E_{\phi}(X)\right\}$ if $\phi$ is admissible. This also follows from Koning’s result after only slightly more work.
@@ 5.1 $\mathbf{U}$-admissibility
The invocation of Condition (C1) in Theorem 5.1 prohibits us from translating the results into the standard Neyman-Pearson setting. That is, Observation 2.6 does not apply. Indeed, for a singleton $\mathcal{B}=\{b\}$, (C1) implies that $L_{b}(1,0)=0$, so type-II risk loses its meaning and does not translate into the power of a test in the traditional paradigm.
@@
Intriguingly, Theorem 5.1 and Corollary 5.3 recover the result of grunwald2024beyond up to some assumptions, even though he studied a weaker form of admissibility. In particular, Theorem 1 of grunwald2024beyond essentially shows that any G-admissible rule is defined by a sharp e-variable. Roughly speaking, he considers a test family to be G-admissible if it is not dominated with probability 1 by another test. (We use G-admissibility to refer to Grünwald.)
@@ 5.2 Improving tests via Rao–Blackwellization
These examples might make us wonder whether all “good” e-variables for post-hoc hypothesis testing for point hypotheses are functions of the likelihood ratio. In general, of course, there are e-variables which are not functions of $\Lambda$. Indeed, $E=\text{d}R/\text{d}P$ for any distribution $R$ is a (sharp) e-variable. However, unless $R=Q$, such an e-variable is not log-optimal (shafer2021testing). We might therefore expect that it can be improved as a post-hoc test.
Here we show that Rao-Blackwellization (radhakrishna1945information; blackwell1947conditional) can weakly improve the type-II risk of any test family based on an e-variable, with respect to a particular set of adversaries. Recall that the Rao-Blackwell theorem says that the expected loss of an estimator can be weakly improved (with respect to a convex loss) by conditioning on a sufficient statistic. See wasserman2004all for a modern statement of the result.
@@ 5.3 $\mathbf{C}$-admissibility
We begin by proving that any randomized $\mathbf{C}$-admissible test must be increasing in the likelihood ratio. The result is, of course, satisfyingly intuitive—a larger likelihood ratio indicates indicates more evidence against the null. A test that rejects the null at $\Lambda(X)$ but fails to reject when $\Lambda(Y)>\Lambda(X)$ is doing so despite increased odds in favor of the alternative. That such a result does not hold for richer classes of adversaries $\Gamma$ is yet another reason to consider $\mathbf{C}$. Indeed, tests which don’t satisfy this property are flaunting the law of likelihood (royall2017statistical).
@@
The proof may be found in Appendix A.13. The idea is simple: Any test which is not increasing in $\Lambda$ can be perturbed such that it becomes strictly better. The details are intricate, however, requiring a careful analysis and strictly more edge cases than is healthy. This result mirrors a result by koning2024continuous (see Theorem 1) with a similar flavor: he shows that e-variables which maximize a fixed expected concave utility under a (simple) alternative must also be increasing in the likelihood ratio. Connecting his result with Lemma 5.7 may be an interesting avenue for future research.
@@
It is worth noting that the properties given in Lemmas 5.7 and 5.8 are both on display for the likelihood ratio test. For continuous $P$ and $Q$, its associated e-value is $E^{\textsc{np}}(X)=L_{b^{*}}(0,1)$ if $\Lambda(X)\geq\kappa(b^{*})$ and 0 otherwise (see Section 3), which is clearly compatible with $\mathcal{L}$. However, in the discrete case, the e-value obeys $E^{\textsc{np}}(X)=L_{b^{*}}(0,1)(1+\gamma)\mathbf{1}\left\{\Lambda(X)\geq\kappa(b^{*})\right\}$, which may not belong to $\mathcal{L}$. Hence, the assumption of continuity in Lemma 5.8 cannot be removed.
We leave it as an open question whether the conditions specified by Lemmas 5.7 and 5.8 (in addition, of course, to the canonical representation in Lemma 4.5) are also sufficient to specify $\mathbf{C}$-admissible tests. We expect the answer is no. We do, however, conjecture that in addition to our extensions of the likelihood ratio test, mixtures of likelihood ratio tests are $\mathbf{C}$-admissible, subject to compatibility. More formally, if $\phi^{\textsc{np}}_{k}$ is the post-hoc likelihood ratio test on loss $b_{k}$ (i.e., the test defined by (3.2)) then we suspect that the test defined by the (sharp) e-variable
@@ 5.3.1 Binary test families
is $\mathbf{C}$-admissible. Indeed, $\Lambda$ is sharp, compatible with $\mathbb{R}_{\geq 0}$ and obviously increasing in itself. This test is satisfying in that it accords well with calls to report the likelihood ratio instead of the p-value (royall1986effect; royall2000probability; berger1987testing; perneger2001sifting). This nicely bridges the Fisherian and Neyman-Pearson perspectives on hypothesis testing: If a study summarizes its findings via $\Lambda(X)$ as a measure of evidence, downstream analysts can reject at whatever level they like, maintaining type-I risk safety.
@@ 6 Summary
Different families $\Gamma$ result in different classes of admissible tests. For point hypotheses, when $\Gamma$ is all mappings from the data to losses we show that all admissible tests correspond to sharp e-variables. When $\Gamma$ is the set of constant mappings and the tests are binary, we give a classification of all admissible tests which recovers the Neyman-Pearson lemma (neyman1933ix).
This work suggests several open questions. For one, in the case of randomized tests, the set of necessary and sufficient conditions for $\mathbf{C}$-admissibility do not match. Can we close the gap and give a classification of all admissible tests in this case? As discussed in Section 5.3, we expect that mixtures of likelihood ratio tests are admissible but we have been unable to prove it. Second, while we have focused mainly on point hypotheses, hypothesis testing in practice typically involves composite hypotheses. While some of our results extend straightforwardly to composite settings, [^4] In particular, Lemmas 4.1, 4.2, and 4.3 can be extended to composite settings if handled correctly. For instance, a composite version of Lemma 4.1 would posit that $\phi(X,b)\geq\delta(X,b)$ $P$-almost surely for all $P\in\mathcal{P}$ and that there exists some $Q\in\mathcal{Q}$ such that $Q(\phi(X,b)>\delta(X,b))>0$. The representations in Lemmas 4.5 and 4.7 likewise hold in composite settings. can we obtain composite versions of our main classification theorems? Third, we have adopted the typical hypothesis testing setup of two actions: either sustaining or rejecting the null. grunwald2024beyond, however, studies multiple actions. How must our results be modified in this case?
@@ Acknowledgments
We thank the anonymous referees for helpful comments which improved the manuscript and Ruodo Wang for pointers towards measurable selection theorems. BC was supported in part by the CWI internship program and in part by NSERC-PGSD, grant no. 567944. BC and AR acknowledge support from NSF grants IIS-2229881 and DMS-2310718. PG was co-funded by the European Union (ERC Advanced Grant FLEX 101142168).
@@ Appendix B On the definition of admissibility
First let us discuss the notion of admissibility used by grunwald2024beyond, which we henceforth refer to as G-admissibility. We say a test family $\delta$ is G-inadmissible if there exists a type-I risk safe test family $\phi$ such that, for all $b\in\mathcal{B}$ and all $Q\in\mathcal{Q}$,
@@
In other words, $\phi$ has loss at most that of $\delta$ with $Q$-probability 1 on all losses, and has lower loss than $\delta$ with positive probability. If no such $\phi$ exists then we say that $\delta$ is G-admissible. We note that grunwald2024beyond defines G-admissibility in terms of type-I losses instead of type-II, after preprocessing the losses to remove any such that $L_{b}(1,a)>L_{b}(1,a)$ and $L_{b}(0,a)>L_{b}(0,a)$ for some $a$, which can be done without loss of generality.
@@
Let us now discuss some relationships between $\Gamma$-admissibility and G-admissibility. First, $\mathbf{C}$-admissibility implies G-admissibility. This is intuitive: If a test family $\delta$ cannot be beat in expectation for fixed $b$, then it cannot be beat with probability 1 either.
Lemma B.1.
$\mathbf{C}$-admissibility implies G-admissibility.
Proof.
We prove the contrapositive. Suppose that $\delta$ is G-inadmissible. So there exists some $\phi$ such that, for all $b\in\mathcal{B}$,
$\phi(X,b)\geq\delta(X,b)\quad Q\text{-almost everywhere for all }Q\in\mathcal{Q},$
and $\phi(X,b_{0})>\delta(X,b_{0})$ for some $b_{0}$ and all $X\in A$ with $Q_{0}(A)>0$ for some $Q_{0}$. Therefore, $\mathbb{E}_{Q_{0}}[\delta(X,b_{0})\mathbf{1}\left\{A\right\}]<\mathbb{E}_{Q_{0}}[\phi(X,b_{0})\mathbf{1}\left\{A\right\}]$ and $\mathbb{E}_{Q_{0}}[\delta(X,b_{0})\mathbf{1}\left\{A^{c}\right\}]\leq\mathbb{E}_{Q_{0}}[\phi(X,b_{0})\mathbf{1}\left\{A^{c}\right\}]$. Thus $\mathbb{E}_{Q_{0}}[\delta(X,b_{0})]<\mathbb{E}_{Q_{0}}[\delta(X,b_{0})]$. Moreover, since $\delta(X,b)\geq\phi(X,b)$ $\mathcal{Q}$-almost surely for all $b$, we have $\mathbb{E}_{Q}[\delta(X,b)]\leq\mathbb{E}_{Q}[\delta(X,b)]$ for all $b$, hence $\phi$ is strictly preferable to $\delta$ with respect to $\mathbf{U}$. Thus G-inadmissibility implies $\mathbf{C}$-inadmissibility. ∎
Note, however, that this proof does not work if we replace $\mathbf{C}$-admissibility by $\mathbf{U}$-admissibility. Consider the following example: Set $\mathcal{B}=[1,\infty)$ and let $X$ take values on $1,\infty)$. Suppose that $\delta(X,b)>\phi(X,b)$ for all $X\neq b$, and for $X=b$ we have $\delta(X,b)<\phi(X,b)$. Since $X=b$ has measure zero, $\delta$ is strictly preferable to $\phi$ with respect to G-admissibility. But for $B(X)=x$ we have $\mathbb{E}_{Q}[L_{B}(1,\delta(X,B))]=\mathbb{E}_{Q}[L_{X}(1,0)(1-\delta(X,X))]>\mathbb{E}_{Q}[L_{X}(1,0)(1-\phi(X,X))]=\mathbb{E}_{Q}[L_{B}(1,\phi(X,B))]$, so $\delta$ is not preferable to $\phi$ with respect to $\mathbf{U}$-admissibility. This shows that we cannot conclude $\mathbf{U}$-inadmissibility from G-inadmissibility.
Of course, this does not provide a counterexample to the claim that $\mathbf{U}$-admissibility implies G-admissibility per se. But it sheds light on the fact that one would need a proof totally unlike the one above. This makes us doubt that the claim is true.
@@ B.1 Type-I risk as the supremum over adversaries
where $\Gamma_{\Omega}$ represents the projection of a set in $\Omega\times\mathcal{B}$ onto $\Omega$. The set $A_{c}$ is Borel (i.e., lies in $\Sigma_{\mathcal{X}}\otimes\mathcal{B}$) by the measurability of $L_{b}(0,1)\delta(X(\omega),b)$. The projection of a Borel set on a product space is analytic by Suslin’s theorem, hence $\{\omega:g(\omega)>c\}$ is analytic (kechris2012classical, Exercise 14.3). Moreover, every analytic set is universally measurable (kechris2012classical, Theorem 21.10), meaning it is measurable for any finite Borel measure on $\Omega$, and $P$ is finite and Borel by assumption. Finally, if the super-level sets of a real-valued function are measurable, then the function itself measurable. We conclude that $g$ is $\mathcal{F}$-measurable, so the left hand side of (2.3) exists. Now, for any $X$ and $B$, $L_{B(X)}(0,\delta(X,B(X)))\leq\sup_{b}L_{b}(0,\delta(X,b))$. Taking expectations and then taking the supremum over all maps $B$, we see that the right hand side of (2.3) is at most the left hand side.
@@
which is $\mathcal{F}$-measurable since $(\omega,b)\mapsto L_{b}(0,\delta(X(\omega),b))$ is measurable. The Kuratowski–Ryll-Nardzewski measurable selection theorem (kuratowski1965general) (see also kechris2012classical) guarantees the existence of a measurable selector $B^{*}_{\epsilon}(X(\omega))$ such that $B_{\epsilon}(X(\omega))\in\Xi_{\epsilon}(\omega)$, hence $L_{B^{*}_{\epsilon}(X)}(0,\delta(X(\omega),B_{\epsilon}^{*}(X))+\epsilon>\sup_{b}L_{b}(0,\delta(X,b))$. Taking expectations and letting $\epsilon\downarrow 0$ shows that the left hand side of (2.3) is at most the right, which proves the result. ∎
@@
There exist pathological choices of $\{L_{b}\}_{b\in\mathcal{B}}$ which ensure that $\sup_{b}L_{b}(0,\delta(X,b))$ is not measurable, in which case Definition 2.3 breaks down and Lemma B.2 doesn’t hold. Instead of making various measurability assumptions as we do in this work, one could circumvent this problem by following grunwald2024beyond and calling $\delta$ type-I risk safe if there exists some measurable function $\xi:\mathcal{X}\to\mathbb{R}_{\geq 0}$ such that